Calculate the weight of the following buffers: CAPS, Acetate, Citrate, and HEPES, that is needed to make 100ml of a 0.1M buffer.
Answer #1

The weight of each salt required to prepare the buffer is calculated:

m CAPS = M * V * MM = 0.1 M * 0.1 L * 221.32 g / mol = 2.21 g

m Sodium Acetate = 0.1 * 0.1 * 82 = 0.82 g

m Sodium Citrate = 0.1 * 0.1 * 258 = 2.58 g

m HEPES = 0.1 * 0.1 * 238 = 2.38 g

If you liked the answer, please rate it in a positive way, you would help me a lot, thank you.